You purchased a house 10 years ago at a cost of $285,000. Today, you sold that house for $500,000. What annual rate of return did you earn on this house?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

500,000=285,000*(1+r/100)^10

(500,000/285,000)^(1/10)=(1+r/100)

(1+r/100)=1.0578

r=1.0578-1

=5.78%(Approx).